### v3

# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

Abstract
We investigate whether the pre-trained knowledge of vision-language models (VLMs), such as CLIP, can be retained or even enhanced during continual learning (CL) while absorbing knowledge from a data stream. Existing methods often rely on additional reference data, isolated components for distribution or domain predictions, leading to high training costs, increased inference complexity, and limited improvement potential for pre-trained models. To address these challenges, we first comprehensively analyze the effects of parameter update locations and ranks on downstream adaptation and knowledge retention. Based on these insights, we propose Dynamic Rank-Selective Low Rank Adaptation (LoRA), a universal and efficient CL approach that adaptively assigns ranks to LoRA modules based on their relevance to the current data. Unlike prior methods, our approach continually enhances the pre-trained VLM by retaining both the pre-trained knowledge and the knowledge acquired during CL. Our approach eliminates the need for explicit domain or distribution prediction and additional reference data, enabling seamless integration of new tasks while preserving pre-trained capabilities. It also maintains the original architecture and deployment pipeline of the pre-trained model without incurring any additional inference overhead. Extensive experiments and analyses demonstrate that our method outperforms state-of-the-art approaches in continually absorbing knowledge of downstream tasks while retaining pre-trained knowledge.
1 Introduction
Figure 1: Illustrations of approaches for CL of VLMs. (a) Continual full fine-tuning is applied toward a universal CL updating. Additional reference data is used to alleviate the forgetting of pre-trained knowledge (zero-shot prediction on unseen data), e.g., [62]. It is computationally expensive and suffers from forgetting. (b) New task-specific components isolated to the pre-trained model are added in CL, e.g., [57, 54]. They require task/domain prediction/gating in inference, increasing inference complexity and restricting the pre-trained model’s improvement potential with limited usage scenarios. (c) We aim to continually learn the model by memory-efficient updating of all parameters in a universal manner, similar to (a), but without requiring reference data. It allows for enhancing the pre-trained model while learning continuously from the data stream, surpassing the limitations of (b). By utilizing dynamic rank-selective LoRA, we reduce forgetting of both the knowledge in the pre-trained model and the knowledge acquired during continual learning.
Continual Learning (CL) [16, 9, 47, 25, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [36, 32] or the need to retrain from scratch. With the advances of large-scale pre-trained models (PTMs), recent CL advances are increasingly leveraging these models. While many PTM-based CL approaches address continual adaptation for downstream tasks (e.g., class-incremental learning [51, 46, 20]), with an emphasis on downstream tasks’ performance, there is growing interest in preserving the inherent capabilities of PTMs, such as multi-domain continual learning for Vision-Language Models (VLMs) [62, 57, 54], or in continual pre-training to enhance PTMs over time [41]. In this paper, we focus on multi-domain continual learning for VLMs like CLIP [39].
The ZSCL [62] framework aims to continually improve pre-trained model by incrementally feeding it with new knowledge. However, this often results in a degradation of both pre-trained and previously acquired knowledge during continual learning steps. To address this challenge, prior works have explored strategies like using reference data [62] or memory replay [25, 51, 40] to distill knowledge during full fine-tuning. However, these methods incur high training costs due to the need for additional data and extensive retraining (Fig. 1(a)), and yet they still suffered from the forgetting problem. Other approaches [57, 54] introduce task-specific components isolated from the pre-trained model during the CL stream, which require task/domain prediction or gating at inference. This increases inference complexity and limits the model’s adaptability to unseen data, restricting its generalization potential (Fig. 1(b)).
In this work, we explore a more general and practical form of CL for VLMs (similar to [62]), which maintains strong performance on seen data with continually updated pre-trained model, and retains or even boosting the performance on unseen data. To ensure memory-efficient learning, we utilize low-rank parameter updates (LoRA) [19]. Rather than inserting LoRA modules at fixed locations [26, 15], we extend them to all pre-trained weight matrices to enhance adaptation, and crucially, to analyze the impact of each weight matrix on downstream learning and knowledge retention.
Our analysis (Sec. 3.2) reveals that not all parameter updates contribute equally to downstream adaptation, with some potentially harming pre-trained capabilities. Manually selecting update locations and ranks can significantly improve the balance between adapting to new tasks and retaining pre-trained knowledge. This insight motivates our proposed method, Dynamic Rank-Selective LoRA, which adaptively determines both the locations within the pre-trained weight matrices and the ranks of parameter updates.
We leverage a reformulation of parameter updates using Singular Value Decomposition (SVD) [60, 12, 28, 35, 59]. Our method introduces learnable importance weights that mimic the singular values in SVD, capturing the significance of each rank in the parameter update. These learnable weights enable the model to dynamically adjust the contribution of each rank during training. To ensure only the most important updates are integrated into the pre-trained weights, we optimize the importance weights using a sparse prompting approach with a soft-thresholding operation. Our contributions are fourfold:
  1. 1.
    In this paper, we explore the universal CL of VLMs in an memory-efficient manner, maintaining strong performance on both seen and unseen data with reduced forgetting. We perform dynamic rank selective parameter updates adaptively to all pre-trained weight matrices.
  2. 2.
    Through comprehensive analyses, we show that the location and rank of parameter updates are critical for achieving an good balance between downstream task adaptation and pre-trained knowledge retention.
  3. 3.
    We propose a general low-rank parameter update method that adaptively perform dynamic rank selective updates, effectively integrating new knowledge while maximizing the retention of pre-trained knowledge.
  4. 4.
    Extensive experiments and in-depth analyses validate our approach across a wide range of benchmarks for downstream adaptation and zero-shot transfer. Additionally, we evaluate on hold-out data to further evaluate knowledge retention.
2 Related Works
Continual Learning. Continual Learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods [30, 3, 6, 27, 7, 55, 56] store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches [22, 1, 58, 2, 20] constrain updates to important weights to preserve past knowledge. Meanwhile, dynamic networks [44, 45, 64, 46, 34, 26, 50, 49, 42, 48] adjust their architecture on the fly and balance learning new information while retaining old knowledge.
Multi-Domain Continual Learning of VLMs. Continual learning of VLMs [21, 61] aims to enable VLMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. The approach introduced in ZSCL [62] addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting. Subsequent works [57, 43, 54] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, [54] proposed a more challenging evaluation setting, where the labels from different domains are mixed during testing.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) [19] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [59, 35], treating each rank independently [12, 60, 28], or employing a mixture of ranks [52]. These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
3 Methodology
3.1 Preliminaries
Continual Learning of VLMs. We focus on continual learning for VLMs within two challenging settings: multi-domain task-incremental learning (MTIL) [62, 57] and cross-domain task-agnostic incremental learning (X-TAIL) [54]. In the MTIL setting, the model sequentially learns from a set of $T$ tasks. For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task. The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ consists of the class names within the task $t$, containing a total of $M^{t}$ classes.
In the X-TAIL setting, however, the category set during testing encompasses both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$. Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training. This introduces a more challenging scenario, as the model must classify test images across a diverse range of domains, making it more complex than the MTIL setting.
Transformers. A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multi-layer perceptron (MLP). In each transformer layer, the attention module at the $l$-th transformer layer can be defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation. The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively.
The MLP module within each transformer block consists of two linear transformations separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU [18]. The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension of the MLP.
Vision-Language Models (VLMs). In vision-language models like CLIP [39], the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case, respectively. Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$. This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
Figure 2: Evaluations of adaptation to downstream tasks and retention of zero-shot capabilities after training CLIP with various LoRA insertion locations and ranks.
Figure 3: Analysis of the effects of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
Figure 4: Overview of our proposed method: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities. After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
3.2 Analyses on LoRA Location and Rank
Analyses of locations and ranks of low-rank updates. We begin by investigating the influence of performing parameter updates at different locations within the pre-trained model, varying the update strength to assess its effects on both downstream adaptation and retention of pre-trained capabilities. The results of this analysis are presented in Fig. 3. We leverage Low-Rank Adaptation (LoRA) [19], where low-rank parameter updates are directly applied to the pre-trained weight matrices.
Our findings reveal the following: (1) Updating only the vision encoder leads to catastrophic forgetting of the pre-trained capabilities, significantly diminishing zero-shot performance. (2) Updating only the attention modules generally reduces forgetting but restricts the model’s ability to adapt to new tasks. (3) Using a higher rank enhances adaptation to downstream tasks without necessarily increasing the forgetting of pre-trained knowledge.
These observations highlight the need for a strategic selection of insertion locations and ranks to achieve an better balance between adaptation and zero-shot retention.
Analyses of pruning subsets of low-rank updates. To further investigate, we analyzed the effects of selectively perturbing individual modules of the pre-trained weights, as shown in Fig. 3. Initially, we inserted LoRA into all pre-trained matrices with a rank of 16, corresponding to the “Vision + Text, Attn + MLP, rank 16” configuration shown in Fig. 3. We then iteratively remove the LoRA updates learned for each module, examining the impact on both downstream adaptation and the knowledge retention.
Our findings reveal that not all parameter updates are equally critical for downstream adaptation, and some may even degrade the model’s pre-trained capabilities. For example, removing updates in the Value projection of the Attention module within the vision encoder had minimal impact on downstream adaptation but significantly restored the model’s pre-trained zero-shot performance. Similarly, removing updates in the MLP layer of the vision encoder slightly reduced downstream task performance but, interestingly, enhanced the model’s pre-trained capabilities, even surpassing its original zero-shot performance.
3.3 CL VLMs with Dynamic Rank-Selective LoRA
Overview. As shown in Fig. 4, we introduce dynamic rank-selective parameter updates for all pre-trained weight matrices, adaptively inserting updates based on the importance of each rank for current training data of task $t$. Specifically, we integrate dynamic rank-selective LoRA into all pre-trained weights within the Attention and MLP modules of both the vision and text encoders. For clarity, for learning on task $t$, we denote the pre-trained weights or those updated from the previous task as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices added with LoRA. After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
    Method Aircraft [31] Caltech101 [14] DTD [8] EuroSAT [17] Flowers [37] Food [5] MNIST [11] OxfordPet [38] Cars [23] SUN397 [53] Average
CLIP
   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
Transfer
   Zero-shot [39] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF [25] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT [51] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL [40] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL [62] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
[…]
   Ours 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
[…]
Average
   LwF [25] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT [51] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL [40] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL [62] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
[…]
   Ours 41.4 81.0 58.6 77.8 83.4 84.5 64.5 90.3 67.2 64.4 71.3
[…]
Last
   LwF [25] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT [51] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL [40] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL [62] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
[…]
   Ours 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
[…]
Table 1: Comparison of different CL methods on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [57, 54].
Rank-selective LoRA with importance weights. LoRA [19] is a parameter-efficient fine-tuning technique designed to adapt pre-trained models using a minimal number of additional parameters. Instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is defined as:
$\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\Delta\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\mathbf{B}^{t,m}\mathbf{A}^{t,m}.$ (1)
In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only the matrices $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, significantly reducing the number of trainable parameters to $r(d+k)$. Notably, during inference, LoRA incurs no additional computational overhead compared to the original pre-trained model, as the low-rank updates can be merged into the pre-trained weights, as shown in Eq. (1).
Previous works reformulate the low-rank parameter update in the form of Singular Value Decomposition (SVD) [60, 12, 28, 35, 59]. We introduce dynamic importance weights into the formulation:
$\Delta\mathbf{W}^{t,m}=\sum_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
where $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ is a learnable vector that represents the importance of each rank during training. The importance weights imitate the diagonal elements in the singular value matrix of SVD and we allow these weights to be learned directly from the data.
The introduction of these importance weights incurs only a minimal increase in trainable parameters, adding just $r$ additional parameters per LoRA module. For initialization, we adhere to the convention that randomly initialize the matrix $\mathbf{A}$, set the matrix $\mathbf{B}$ to zero, and we initialize the importance weights $\mathbf{w}^{t,m}$ randomly.
Dynamic rank selection with sparse importance. The formulation in Eq. (2) enables the model to dynamically adjust the importance of each rank through gradient descent. [^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss during pre-training as the supervised training loss. However, not all ranks contribute equally to downstream task performance (as shown in Fig. 3). To ensure that only the most impactful LoRA ranks are retained in the pre-trained weights, we propose to optimize the learning of importance weights in a sparse-promoting way by minimizing the $\ell_{1}$ norm of the importance weights. The optimization objective for paramters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ are as follows:
$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ denotes the total number of pre-trained weight matrices inserted with LoRA of rank $r$, and $\mathbf{w}^{t,m}$ represents the importance weights of the each rank of the LoRA module added to the pre-trained weight matrix $\mathbf{W}^{t,m}_{0}$, and $\lambda$ controls the strength of the $\ell_{1}$ regularization.
We adopt the proximal gradient [4] for handling the non-differentiable $\ell_{1}$ regularization applied to the importance weights. Specifically, the dynamic update of the importance weight for the $i$-th rank of LoRA can be written as a soft-thresholding operation:
$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}+\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after gradient update using only the supervised loss $\mathcal{L}_{\text{sup}}^{t}$. The threshold parameter $\kappa$ is initially set to zero and gradually increases until it reaches a maximum value $\kappa_{\text{max}}$. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if condition is met, 0 otherwise, and $\mathrm{sign}(\cdot)$ represent the +/+/-+ / - sign of input.
This encourages that only the ranks with significant importance are preserved, while those with low importance are pruned. By adjusting the weightings iteratively, the model focuses dynamic parameter updates on the most relevant components. In the early stages of training, dense updates are performed without Eq. (4) to allow all ranks to capture task-relevant information.
Continually enhance the pre-trained model. After training is complete, ranks with importance weight of zero are removed, while non-zero ranks are merged into the pre-trained weights by substituting Eq. (2) into the original LoRA formulation Eq. (1). When a new task is introduced, a new set of LoRA modules is initialized, and the computations involve only the newly initialized LoRA modules in conjunction with the continually updated pre-trained model. This approach allows the model to leverage accumulated knowledge efficiently while keeping the computational footprint minimal without additional overhead.
    Method Aircraft [31] Caltech101 [14] CIFAR100 [24] DTD [8] EuroSAT [17] Flowers [37] Food [5] MNIST [11] OxfordPet [38] Cars [23] SUN397 [53] Average
CLIP
   Zero-shot [39] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
Transfer
   Zero-shot [39] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF [25] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR [13] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT [51] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL [62] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
[…]
   Ours 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
[…]
Average
   LwF [25] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR [13] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT [51] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL [62] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
[…]
   Ours 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
[…]
Last
   LwF [25] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR [13] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT [51] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL [62] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
[…]
   Ours 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
[…]
Table 2: Comparison with state-of-the-art methods on 5-shot MTIL setting in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [57, 54].
4 Experiments
4.1 Experimental Setting
Datasets. We evaluate our method on two widely used continual learning settings for vision-language models: MTIL [62, 57] and X-TAIL [54]. For the MTIL setting, a total of 11 datasets are used, where each dataset is treated as an individual task to be incrementally learned. For the X-TAIL setting, CIFAR100 is excluded to prevent domain overlap, following the protocol in [54]. In line with [54], we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
Evaluation Metrics. We adopt the same evaluation metrics used in prior works, which include “Transfer”, “Average”, and “Last”. The “Transfer” metric evaluates the model’s zero-shot transfer capability on previously unseen data. The “Last” metric assesses the model’s ability to retain knowledge from earlier tasks. The “Average” metric is a composite score that reflects the mean performance across both the “Transfer” and “Last” evaluations.
Implementation Details. We follow the experimental setups in [62, 57, 54] and utilize the CLIP model with a ViT-B/16 backbone [39] for all experiments. By default, our proposed importance-aware sparse LoRA is applied to all pre-trained weight matrices across both the vision encoder and text encoder, with an initial rank set to 16. Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [29]. The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations. The maximum threshold value for adaptive pruning is fixed at 0.005 for both settings.
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [39] 68.24 33.77 66.72 56.24
MoE-Adapter [57] 68.24 33.77 66.72 56.24
RAIL [54] 68.24 33.77 66.72 56.24
Ours 68.95 36.52 68.52 58.00
Table 3: Evaluation of average transfer accuracy on datasets that were never encountered during the continual learning stages.
Methods
Training
Params.
Additional Infer.
Params. / Mem.
LWF [25] 129.6M None
ZSCL [62] 129.6M None
MoE-Adapters [57] 59.8M 13.35M
RAIL [54] N/A 24.18M / 9.01M
Ours 4.4M None
Table 4: Comparison on computation cost with previous methods.
4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report the evaluation results on the X-TAIL setting in Table 1. Unlike previous works that rely on model parameter updates [25, 51, 40, 13] along with additional memory replay or reference data, our approach performs parameter-efficient updates without requiring any memory replay or reference data. This results in significant improvements across all evaluated metrics.
In contrast to prior methods [57, 54] [^2] Dual version of [54] stores all the training samples’ feature embeddings in the memory, for a fair comparison, we excluded the results here. We leave more detailed comparisons in Appendix., which use zero-shot predictions to preserve pre-trained capabilities and prevent forgetting, our method overcomes this limitation by continually refining the model. Notably, while these approaches are constrained by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy. Although some predictions on unseen data might be slightly impacted due to the nature of continual parameter updates, our method ultimately achieves a higher overall transfer accuracy by effectively adapting the pre-trained model to new tasks while also enhancing its generalization capabilities.
For both the Average and Last metrics, our method demonstrates superior or comparable performance without relying on domain ID prediction [57], high-dimensional projections, or storing a memory bank of previously seen sample features [54]. Instead, our method focuses purely on updating the pre-trained model, yet still delivers competitive results.
Furthermore, since our approach is entirely orthogonal to the additional techniques used in previous methods [57, 54], these techniques can be seamlessly integrated into our framework to further enhance performance, benefited from our ultimate goal of continually enhancing the pre-trained model’s capabilities.
Multi-domain task-incremental learning. We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [57, 54]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this setting. Without employing any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average Transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works. Moreover, by supplementing our method with techniques used in prior approaches [57, 54], we can further improve the performance.
4.3 Discussions
Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset using our proposed method.
Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset using our proposed method.
Pre-trained knowledge retention and potentially improved pre-trained model. To understand how our method improves transfer performance beyond the upperbound set by zero-shot predictions in previous works, we evaluate on datasets that were never encountered during the continual learning steps, specifically CIFAR100 [24], Places365 [63], and ImageNet-1k [10], as shown in Table 3.
Unlike prior approaches [57, 54] that rely exclusively on zero-shot predictions for unseen data, thereby limiting their performance to that of the original CLIP model, our method goes beyond this limitation. By effectively integrating new knowledge into the pre-trained parameter space without sacrificing existing capabilities, our approach enhances the model’s overall performance. Consequently, our method achieves superior transfer accuracy, surpassing the theoretical upper bound defined by zero-shot performance in previous methods. This not only demonstrates improved adaptation but also enhances the model’s generalization capabilities on completely unseen datasets during CL.
Rank allocation results on different datasets. We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig. 6 and Fig. 6. Overall, it is evident that the Aircraft dataset generally requires more ranks to be updated in the pre-trained model compared to the Pets dataset.
We observe that for the Aircraft dataset, a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig. 6(b)). Additionally, a higher concentration of ranks is assigned to the deeper layers, particularly layers 6 to 10 (Fig. 6(c)).
In contrast, the analysis of the Pets dataset (Fig. 6(b) and Fig. 6(c)) reveals that more ranks are allocated to the Attentions’ Key and Value projection, and MLP’s FC layer of vision encoder. Furthermore, the majority of rank allocations are in the first and last layers, suggesting a different pattern for promoting knowledge retention and adaptation across datasets. These findings highlight the adaptive nature of our approach, where rank allocation varies based on the characteristics of the dataset, optimizing the model’s ability to capture domain-specific information effectively.
Computation Cost. We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig. 4. For MoE-Adapter, the difference between training and testing parameters stems from its mixture-of-experts design: during training, all experts are active, while at inference, only the top-2 experts are selected, reducing the active parameter count. RAIL-Primal incurs additional computational overhead during inference due to its use of high-dimensional feature expansion through projection and regression layers. RAIL-Dual maintains a memory bank of image features for all training samples across 1,100 classes, which significantly increases memory usage.
Notably, these methods experience linearly growing computational costs as the number of tasks increases. In contrast, our method minimizes training overhead by selectively retaining only ranks with high importance. Once training is complete, these updates are merged into the pre-trained weights, resulting in zero additional memory overhead during inference.
Figure 7: Ablation studies on different insertion locations and the impact of varying the initial rank.
Figure 8: Analyses of the effects of (a) varying the maximum pruning threshold and (b) different ratios of dense training iterations.
4.4 Ablation Studies
Effects of insertion locations and number of initial ranks. We examine the impact of different insertion locations and initial ranks for our method, with results presented in Fig. 7. These findings are consistent with our analytical experiments shown in Fig. 3. Our results indicate that when training is restricted to only the vision encoder, transfer accuracy is significantly reduced, suggesting that the pre-trained capabilities are negatively impacted. Conversely, increasing the maximum rank improves adaptation ability without necessarily reducing transfer accuracy. Furthermore, when limiting insertion to only the Attention modules, we observe that transfer accuracy can surpass our previously reported state-of-the-art results in Table 1 and Table 2. However, this comes at the expense of reduced adaptation capabilities. To achieve an better balance between adaptation and retention, we enable insertion into all modules by default. This configuration ensures a better trade-off between adapting to new tasks while preserving the model’s zero-shot transfer capabilities.
Effect of maximum threshold value and ratio for dense iterations. We analyze the effects of varying the maximum threshold value $\kappa_{\text{max}}$ and the ratio for dense training iterations in Fig. 8. The results indicate that as the threshold increases, the Transfer accuracy shows an increasing trend. This suggests that pruning more ranks can enhance the model’s performance on unseen data (i.e., improve transfer capability), but it also limits the model’s ability to adapt to downstream tasks, as evidenced by changes in the Last accuracy metric. Our experiments show that a threshold value around 0.005 strikes a good balance between retaining pre-trained knowledge and adapting to new tasks.
The ratio of dense training iterations determines when the model begins the adaptive pruning process. Setting this ratio to 0 (i.e., starting pruning from the beginning of training) severely hinders adaptation, as LoRA lacks sufficient time to learn essential task-related knowledge before pruning begins. On the other hand, delaying pruning for too long can result in the model forgetting pre-trained capabilities, thereby reducing transfer accuracy. Our findings indicate that setting the ratio around 0.5 to 0.7 works best across all experimental settings, providing a good balance between adaptation and knowledge retention.
5 Conclusion
In this paper, we introduced a novel Dynamic Rank-Selective LoRA approach to address the challenges of continual learning for vision-language models. By dynamically adjusting the ranks based on each ranks’ importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead. Extensive evaluations on multi-domain and cross-domain settings demonstrated that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and zero-shot generalization.
Limitations and future works. Our current approach directly merges parameter updates into the pre-trained weights without explicitly considering correlations between tasks. Exploring methods to leverage inter-task relationships and shared information build a promising direction for future research.
Acknowledgment
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
Supplementary Material
Appendix A Additional Details of the Proposed Method
In Sec. 3.3 of the main paper, we introduce our method for learning dynamic rank-selective parameter updates to pre-trained weight matrices. We leverage a learnable vector to represent the importance of each rank, which is updated through a soft-thresholding operation. In this section, we provide more details to enhance clarity and understanding.
Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper.
A.1 Full Derivation of Eq. (4)
For simplicity, we omit the notation for the training task $t$ in the following derivations. With a slight abuse of notation, we use subscript $t$ to represent $t$-th training iteration. At the $t$-th training iteration, the training loss with $\ell_{1}$ sparse regularization, as defined in the main paper, is:
$\mathcal{L}_{\text{train}}(\Delta_{t}):=\mathcal{L}_{\text{sup}}(\Delta_{t})+\lambda\sum_{m=1}^{M}\|\mathbf{w}^{m}_{t}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}$ is the supervised training loss, $\Delta_{t}$ is the all trainable parameters $\{\mathbf{w}^{m}_{t},\mathbf{B}^{m}_{t},\mathbf{A}^{m}_{t}\}_{m=1}^{M}$ at $t$-th iteration, $M$ is the total number of pre-trained weight matrices with LoRA of rank $r$, $\mathbf{w}^{m}_{t}$ represents the importance weights of the $m$-th LoRA module, and $\lambda>0$ controls the strength of the $\ell_{1}$ regularization.
We apply proximal gradient descent [4] to handle the non-differentiable $\ell_{1}$ regularization of the importance weights [12]. At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
$\mathbf{w}_{t+1}^{m}$ $\leftarrow\arg\min_{\mathbf{w}^{m}}\,\eta_{t}\cdot\lambda\|\mathbf{w}^{m}\|_{1}$
$\quad+\frac{1}{2}\|\mathbf{w}^{m}-(\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t}))\|_{2}^{2},$ (5)
where $\eta_{t}>0$ is the step-size of $t$-th iteration, and $\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})$ is the gradient of the supervised loss with respect to $\mathbf{w}^{m}$. This update balances the $\ell_{1}$ sparsity regularization term $\|\mathbf{w}^{m}\|_{1}$, with the quadratic penalty term $\frac{1}{2}\|\mathbf{w}^{m}-\hat{\mathbf{w}}^{m}_{t}\|_{2}^{2}$ , which ensures proximity to the gradient update if importance weights $\hat{\mathbf{w}}^{m}_{t}:=\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})$. This can be achieved through a soft-thresholding operator $\mathcal{T}$:
$\mathbf{w}_{t+1}^{m}\leftarrow\mathcal{T}_{\eta_{t}\cdot\lambda}(\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})),$ (6)
where the soft-thresholding operator $\mathcal{T}$ is defined as:
$\mathcal{T}_{\kappa}(x)=\mathbbm{1}(\lvert x\rvert>\kappa)\cdot(x+\mathrm{sign}(x)\cdot\kappa),$ (7)
such that the input to the operator $\mathcal{T}_{\kappa}$ is the elements in $\hat{\mathbf{w}}^{m}_{t}$, $\hat{\mathbf{w}}^{m}_{t}$ is the value of $\mathbf{w}^{m}_{t}$ after the gradient update using only $\mathcal{L}_{\text{sup}}(\Delta_{t})$, and $\kappa=\eta_{t}\cdot\lambda$ is the threshold. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met, and 0 otherwise, while $\mathrm{sign}(\cdot)$ denotes the sign (+/)(+/-)( + / - ) of the input. This concludes our derivation of the dynamic update for the importance weights as presented in the main paper Eq. (4).
Appendix B More Experimental Results and Ablations
B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
Related continual learning methods focus on updating representations through the backbone model (e.g., [50, 62, 57], and ours) or by adding a representation alignment or adaptive prediction head [33, 54]. These two approaches can be considered orthogonal. We investigate how the representation updates introduced by our proposed method can complement the additional adaptive prediction head, utilizing the aggression-based adapters described in [54].
We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 5. Note that the Dual version of RAIL [54] stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 2 for a fair comparison.
Since RAIL heavily relies on the frozen pre-trained model, the Transfer accuracy is limited by the zero-shot performance of the pre-trained model. In contrast, our approach gradually accumulates knowledge into the pre-trained model, enhancing its capability on unseen data and achieving improved Transfer accuracy. Moreover, by leveraging our continually enhanced pre-trained model, coupling our method with RAIL significantly outperforms the original RAIL, which uses a frozen pre-trained model, on the Average and Last metrics.
Transfer Average Last
RAIL-Primal 62.4 70.7 79.1
Ours + RAIL-Primal 63.1 74.1 83.4
RAIL-Dual 62.4 71.9 82.4
Ours + RAIL-Dual 63.1 74.2 83.6
Table 5: Experimental results for integrating aggression-based adapters in the X-TAIL setting. Metrics for “Transfer”, “Average”, and “Last” are reported as averages across all datasets.
B.2 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
In Sec. 4.3 of the main paper, Fig. 6 and Fig. 6 visualize and analyze the dynamic ranks assigned to each module. Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in [19]. The amplification factor is given by:
$\text{Amp}=\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{T}\mathbf{W}\mathbf{V}^{T}\|_{F}},$ (8)
where $\mathbf{W}$ is the pre-trained weight matrix, and $\Delta\mathbf{W}$ is the parameter change introduced by LoRA. For vanilla LoRA and our method, $\Delta\mathbf{W}$ is computed using Eq. (1) and Eq. (2), respectively. $\mathbf{U}$ and $\mathbf{V}$ represent the top $r$ singular vectors of $\text{SVD}(\Delta\mathbf{W})$, where $r$ is the number of interested ranks in analysis.
A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights [19]. We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig. 9.
The results show that, compared to vanilla LoRA, our method tends to concentrate parameter updates on a more specific subset of transformer modules while reducing updates to pre-trained weights less relevant to the current data. This demonstrates that our method more effectively identifies critical pre-trained weight matrices relevant to the current data and applies targeted dynamic parameter updates. Additionally, by reducing the strength of parameter updates for pre-trained weights less related to the current data, our method minimizes the impact on pre-trained knowledge and knowledge acquired from previous tasks compared to vanilla LoRA.
Note that this analysis differs from the visualizations of number of activated ranks in Fig. 6 and Fig. 6. The number of activated ranks reflects the quantity of significant ranks contributing to parameter updates for leaning each task. In this analysis, the amplification factor directly measures the parameter change to each pre-trained weight matrix.
(a) Aircraft - LoRA
(b) EuroSAT - LoRA
(c) Pets - LoRA
(d) Aircraft - Ours
(e) EuroSAT - Ours
(f) Pets - Ours
Figure 9: Visualization of amplification factors when trained on (a, d) Aircraft, (b, e) EuroSAT, and (c, f) Pets, comparing (d, e, f) our method of dynamic sparse rank selective LoRA and (a, b, c) vanilla LoRA. It shows that the proposed method with adaptive rank selection can achieve more focused/concentrated updating and less interference.
B.3 More Visualizations and Analyses of Rank Allocation Across Datasets
In Sec. 4.3, Fig. 6 and Fig. 6, we visualize and analyze the dynamic ranks assigned to each module for the Aircraft and Oxford Pets datasets under the X-TAIL experimental settings. Here, we extend these visualizations and analyses to additional datasets, as shown from Fig. 11 to Fig. 17.
The visualizations, along with statistical analyses of transformer modules and layers, reveal distinct rank allocation patterns across datasets. These findings suggest that the rank of parameter updates needed for achieving a good downstream improvements and knowledge retention is distinct across each kind of data.
Figure 10: Visualization and statistical analysis of rank activation on the Caltech101 dataset using our proposed method.
Figure 11: Visualization and statistical analysis of rank activation on the DTD dataset using our proposed method.
Figure 12: Visualization and statistical analysis of rank activation on the EuroSAT dataset using our proposed method.
Figure 13: Visualization and statistical analysis of rank activation on the Flowers dataset using our proposed method.
Figure 14: Visualization and statistical analysis of rank activation on the Food101 dataset using our proposed method.
Figure 15: Visualization and statistical analysis of rank activation on the MNIST dataset using our proposed method.
Figure 16: Visualization and statistical analysis of rank activation on the Stanford Cars dataset using our proposed method.
Figure 17: Visualization and statistical analysis of rank activation on the SUN397 dataset using our proposed method.