### v5

# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

Abstract
Continual learning (CL) aims to accumulate knowledge from sequential data and task streams. Leveraging their strong generalization and flexibility, pre-trained vision-language embedding models such as CLIP (Contrastive Language-Image Pre-training) have been widely adopted and validated in CL. In addition to learning new knowledge, we investigate whether the pre-trained knowledge in CLIP, can be retained or even enhanced, in CL, while incorporating new knowledge from a data stream. Existing CL methods primarily focus on continual downstream adaptation using components isolated from the pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge by relying on additional reference data, resulting in high training costs. To address these limitations, we propose a universal and efficient CL approach for CLIP based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL. By analyzing how LoRA rank and placement affect learning and forgetting in CL, we design CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data. This ensures a balance between knowledge acquisition (plasticity) and forgetting mitigation (stability). Our method operates without explicit domain or distribution prediction and does not rely on reference data, enabling seamless task integration while maintaining pre-trained capabilities. Moreover, CoDyRA preserves the original model architecture and deployment pipeline, introducing no additional inference overhead. Extensive experiments show that our approach enhances representations for new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
1 Introduction
Continual learning (CL) [18, 9, 49, 29, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [39, 36, 17] or the need to retrain on all previously seen data. With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context [70, 37, 48, 24, 58]. Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL [44, 54, 24, 37]), with an emphasis on downstream task performance. While these task-specific designs [70, 37, 58] yield promising results, preserving the inherent capabilities of PTMs during CL is critical to maintaining and improving generalizable representations. Benefiting from the strong generalization and flexibility of jointly learned vision-text embeddings, vision-language models such as CLIP have been widely adopted and studied in CL [24, 66, 50, 69], including continual updating [17] or multi-domain CL [66] of CLIP.
Some approaches, such as ZSCL [66], aim to update PTM, i.e., CLIP, by incrementally incorporating new data through full fine-tuning while mitigating forgetting via memory replay [29, 54, 43] using additional reference data. However, this leads to high training costs and still struggles with forgetting (Fig. 1(a)). Other methods [61, 58] retain PTM capabilities during downstream CL by introducing task-specific components that are isolated from the PTM, requiring task/domain prediction or gating at inference. This increases inference complexity and limits adaptability to unseen data, restricting the model’s generalization potential (Fig. 1(b)).
Figure 1: CL methods for CLIP: (a) Continual full fine-tuning uses reference data to apply universal updates and mitigate forgetting of pre-trained knowledge (e.g., zero-shot performance on unseen data [66]), but is computationally expensive and still susceptible to forgetting; (b) Task-specific modular methods (e.g., [61, 58]) introduce isolated components that require domain prediction or gating at inference, increasing complexity and limiting adaptability and generalization; (c) Our approach enables universal, parameter- and data-efficient adaptation without reference data or task-specific modules. It continually improves the pre-trained model and achieves superior performance on unseen data.
In this work, we propose a practical and efficient CL approach for CLIP that continually trains and updates pre-trained models without relying on downstream task-specific components. Our method preserves and even enhances the generalizable representations of PTMs while seamlessly integrating new knowledge from data streams. To achieve efficient and effective learning, we conduct comprehensive analyses of low-rank adaptation (LoRA) [21] for CL in PTMs. Rather than restricting LoRA to fixed locations as in prior works [30, 16], we extend it to all parameter weights for a thorough investigation, aiming to approximate the capabilities of full fine-tuning. We analyze how the rank and placement of LoRA updates impact or regularize the learning dynamics, influencing both downstream adaptation and the retention of existing knowledge (Sec. 3.2).
Our analyses reveal that not all LoRA ranks and placements contribute equally to downstream learning and forgetting. Generally, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. However, we identify that there usually exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits. Moreover, we observe that this balance varies significantly across different parameter placements.
The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates by dynamically optimizing for a minimized LoRA rank suitable for the current data. The rank selection for each LoRA is achieved through a Singular Value Decomposition (SVD)-based reformulation [64, 12, 32, 38, 63], with learnable importance weights corresponding to the singular values. With sparsity-promoting regularization, these learnable weights allow the model to dynamically adjust the contribution of each rank during training. This ensures that only the most essential updates (i.e., LoRA with a minimized rank) are integrated into the pre-trained weights, enabling learning from new data while mitigating forgetting. Our contributions are fourfold:
  1. 1.
    We investigate a universal and practical CL paradigm for efficiently updating pre-trained CLIP on new data streams while mitigating forgetting, without reference data or additional components.
  2. 2.
    We conduct in-depth analyses on the impact of LoRA ranks and placements, revealing their crucial role in balancing downstream learning and knowledge retention.
  3. 3.
    We propose CoDyRA, which dynamically optimizes a minimized rank for parameter updates based on current data, with learnable importance weights for each rank.
  4. 4.
    Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, and additionally prove superiority on pre-trained knowledge retention using hold-out data.
2 Related Work
Continual Learning. Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods [34, 3, 6, 31, 7, 59, 60] store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches [25, 1, 62, 2, 23] constrain updates to important weights to preserve past knowledge. Meanwhile, dynamic networks [46, 47, 68, 48, 37, 30, 53, 52, 44, 51] adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
Continual Learning of CLIP. Continual learning of CLIP [24, 65] aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. [17] continually adapt CLIP to temporal distribution shifts via timestamped data updates. The approach introduced in ZSCL [66] addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting. Subsequent works [61, 45, 58] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, [58] proposed a more challenging evaluation setting, where the labels from different domains are mixed during testing.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) [21] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Similar to LoRA, [56] learns a small set of task-specific representation vectors at each layer. Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [63, 38], treating each rank independently [12, 64, 32], or employing a mixture of ranks [55]. These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
Figure 2: Overview of CoDyRA: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities. After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
3 Methodology
3.1 Preliminaries
Transformers. A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multilayer perceptron (MLP). In each transformer layer, the attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation. The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively.
The MLP module within each transformer block consists of two linear transformations separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU [20]. The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension of the MLP.
Vision-language CLIP model. In CLIP model with joint vision-language embedding [42, 22, 14, 28], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case. Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$. This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed using the softmax function as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
CL of CLIP. We focus on CL of CLIP with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [66, 61] and cross-domain task-agnostic incremental learning (X-TAIL) [58]. In the MTIL setting, the model sequentially learns from a set of $T$ tasks. For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task. The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total. In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$. Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training. This poses a more challenging scenario, as the model must classify test images across a broader range of domains.
3.2 Analyses on LoRA Location and Rank
Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention when training CLIP with varying LoRA insertion points and ranks. Shapes indicate encoders, colors denote transformer modules, and sizes reflect rank values.
Figure 4: Analysis of the impact of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
To achieve efficient learning, we explore LoRA and aim to optimize its usage strategy. To analyze the impact of LoRA’s rank and placement on downstream learning and forgetting, we conduct a series of experiments: (1) We apply LoRA components with varying ranks at different locations in the pre-trained CLIP model (Fig. 4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation (Fig. 4). We train the model on downstream datasets and assess its performance on both (1) the new task and (2) an additional reference dataset to evaluate the preservation of pre-trained capabilities.
Applying LoRA on all weights. As shown in Fig. 4 and Fig. 4, applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks. These observations align with the distinct roles of different PTM components. Rather than manually selecting LoRA placement, we apply it across all weights and optimize for an adaptive configuration.
The effects of different LoRA ranks on balancing learning and forgetting. From Fig. 4, we can observe that, despite learning low-rank space, a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. This suggests that a lower-rank LoRA produces stronger regularization in maintaining the learned knowledge by sacrificing the learning of new knowledge. Importantly, we identify a balance at a moderately small rank, which reduces forgetting while maintaining adaptability. This implies that minimizing the rank while fitting on new data may find an optimized trade-off for CL.
The balance varies by location and task. The ranks for a good balance vary by placement (Fig. 4), and pruning LoRA at different locations has varying effects (Fig. 4), indicating that not all LoRA components contributed equally for a task. For example, removing updates in the Value projection of the Attention module in the vision encoder minimally affected adaptation but significantly restored pre-trained performance. Interestingly, pruning the MLP layer of the vision encoder slightly reduced task performance but enhanced pre-trained capabilities. These findings motivate our adaptive approach to dynamically determine ranks based on location and task.
3.3 Proposed CoDyRA
Overview. Our analyses in Sec. 3.2 suggest applying LoRA across all CLIP modules while adaptively selecting an appropriate rank for each task and location to balance learning and forgetting. While finding an ideally global optimal rank is challenging, we approach it by jointly optimizing downstream task performance and minimizing the number of ranks using sparsity-promoting regularization.
We introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders, as shown in Fig. 2. These updates are adaptively inserted based on the importance of each rank for the current training data. For clarity, during training on task $t$, we denote the pre-trained weights—or those updated from the previous tasks—as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices augmented with LoRA. After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
Low-rank updates with rank importance. For efficient adaptation, instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA [21] introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is then defined as:
$\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\Delta\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\mathbf{B}^{t,m}\mathbf{A}^{t,m}.$ (1)
In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, reducing the number of trainable parameters to $r(d+k)$. Notably, LoRA incurs no additional computational overhead during inference, as the low-rank updates can be efficiently merged into the pre-trained weights, as shown in Eq. (1).
To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate it into the formulation of LoRA as follows:
$\Delta\mathbf{W}^{t,m}=\sum\nolimits_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
where $\mathbf{w}^{t,m}$ is learnable and represents the importance of each rank. The formulation is based on a reformulation of LoRA updates in Eq. (1) based on Singular Value Decomposition (SVD) [64, 12, 32, 38, 63]. The importance weights are corresponding to the diagonal elements in the singular value matrix of SVD.
Introducing importance weights incurs minimal overhead, adding only $r$ additional parameters per module. For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
Table 1: Comparisons on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [61, 58].
    Method Aircraft [35] Caltech101 [15] DTD [8] EuroSAT [19] Flowers [40] Food [5] MNIST [11] OxfordPet [41] Cars [26] SUN397 [57] Average
CLIP
   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
Transfer
   Zero-shot [42] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF [29] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT [54] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL [43] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL [66] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
[…]
   CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
[…]
Average
   LwF [29] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT [54] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL [43] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL [66] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
[…]
   CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
[…]
Last
   LwF [29] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT [54] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL [43] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL [66] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
[…]
   CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
[…]
Dynamic rank-selective updates with sparse rank importance. The formulation in Eq. (2) enables the model to dynamically adjust the importance of each rank through gradient descent. [^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training. However, not all ranks contribute equally to downstream task performance (as shown in Fig. 4). To ensure that only the most impactful ranks are retained for learning on new task while removing unnecessary ranks to mitigate forgetting of already-learned knowledge, we propose optimizing the importance weights with a sparsity-promoting regularization through minimizing the $\ell_{1}$ norm. The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ is the total number of pre-trained weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$. The $\ell_{1}$ regularization strength is controlled by $\lambda$.
To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method [4]. The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}+\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after applying the gradient update from the supervised loss $\mathcal{L}_{\text{sup}}^{t}$. The threshold parameter $\kappa$ is initially set to zero and gradually increases to a maximum value $\kappa_{\text{max}}$, analogous to linearly scheduling the regularization strength $\lambda$. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met and 0 otherwise, while $\mathrm{sign}(\cdot)$ represents the +/+/-+ / - sign of the input.
This approach preserves only the ranks with significant importance, pruning those with low relevance. By iteratively adjusting weightings, the model focuses dynamic parameter updates on the most relevant components. During the early stages of training, dense updates are performed without Eq. (4) to allow all ranks to capture task-relevant information before applying sparsity constraints.
Continually enhancing the pre-trained model. After training, ranks with zero importance weights are pruned, while non-zero ranks are merged into the pre-trained weights by substituting Eq. (2) into the formulation Eq. (1). When a new task is introduced, a fresh set of CoDyRA modules is initialized, and computations involve only these newly initialized CoDyRA modules alongside the continually updated pre-trained model. This approach enables the model to efficiently leverage accumulated knowledge while maintaining a minimal computational footprint, ensuring no additional overhead.
Table 2: Comparisons on 5-shot MTIL setting, evaluated in a similar manner as in Table 1.
    Method Aircraft [35] Caltech101 [15] CIFAR100 [27] DTD [8] EuroSAT [19] Flowers [40] Food [5] MNIST [11] OxfordPet [41] Cars [26] SUN397 [57] Average
CLIP
   Zero-shot [42] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
Transfer
   Zero-shot [42] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF [29] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR [13] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT [54] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL [66] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
[…]
   CoDyRA 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
[…]
Average
   LwF [29] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR [13] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT [54] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL [66] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
[…]
   CoDyRA 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
[…]
Last
   LwF [29] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR [13] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT [54] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL [66] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
[…]
   CoDyRA 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
[…]
4 Experiments
4.1 Experimental Setting
Datasets. We evaluate our method on two widely used continual learning settings for vision-language models: MTIL [66, 61] and X-TAIL [58]. For the MTIL setting, a total of 11 datasets are used, where each dataset is treated as an individual task to be incrementally learned. For the X-TAIL setting, CIFAR100 is excluded to prevent domain overlap, following the protocol in [58]. In line with [58], we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
Evaluation Metrics. We adopt the same evaluation metrics as prior works, including “Transfer”, “Average” , and “Last”. “Transfer”evaluates the model’s zero-shot transfer capability on unseen data. “Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last”.
Implementation Details. We follow the experimental setups in [66, 61, 58] and use the CLIP model with a ViT-B/16 backbone [42] for all experiments. By default, CoDyRA is applied to all pre-trained weight matrices in both the vision and text encoders, with an initial rank of 16. Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [33]. The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations. The maximum threshold for adaptive pruning $\kappa$ is fixed at 0.005 for both settings.
4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report the X-TAIL setting evaluation results in Table 1. In contrast to prior methods [61, 58] , which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model. Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy. Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
[…]
Multi-domain task-incremental learning. We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [61, 58]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario. Without incorporating any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works. Furthermore, by integrating techniques from previous approaches [61, 58], our method can be further enhanced, leading to even greater performance gains.
4.3 Discussions
Pre-trained knowledge retention and potential pre-trained model improvement. To understand how our method improves transfer performance beyond the upper bound set by zero-shot predictions in prior works, we evaluated on hold-out datasets never encountered during continual learning, specifically CIFAR100 [27], Places365 [67], and ImageNet-1k [10], as shown in Table 3.
Table 3: Average transfer accuracy on unseen datasets not encountered during CL.
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [42] 68.24 33.77 66.72 56.24
MoE-Adapter [61] 68.24 33.77 66.72 56.24
RAIL [58] 68.24 33.77 66.72 56.24
CoDyRA 68.95 36.52 68.52 58.00
Unlike prior approaches [61, 58], which rely exclusively on zero-shot predictions for unseen data—thereby constraining their performance to that of the original CLIP model—our method overcomes this limitation. By effectively integrating new knowledge into the pre-trained parameter space while preserving existing capabilities, our approach enhances overall model performance. As a result, it achieves superior zero-shot transfer accuracy, surpassing the zero-shot performance limits of prior works. This demonstrates that CoDyRA effectively incorporates general knowledge into its knowledge base, further improving its overall representation capability.
Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset.
Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset.
Rank allocation across different datasets. We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig. 6 and Fig. 6. Overall, the Aircraft dataset generally requires more rank updates in the PTM than the Pets dataset.
Table 4: Computation cost.
Methods
Training
Params.
Additional Infer.
Params. / Mem.
LWF [29] 129.6M None
ZSCL [66] 129.6M None
MoE-Adapters [61] 59.8M 13.35M
RAIL [58] N/A 24.18M / 9.01M
CoDyRA 4.4M None
For the Aircraft dataset, we observe that a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig. 6(b)). Additionally, a higher concentration of ranks is assigned to deeper layers, particularly layers 6 to 10 (Fig. 6(c)).
In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig. 6(b)). Additionally, rank allocation is primarily concentrated in the first and last layers (Fig. 6(c)), indicating a distinct adaptation pattern that balances knowledge retention and adaptation across datasets. These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge. We provide more visualizations and analyses in Appendix.
Computation cost. We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig. 4. For MoE-Adapter, the difference between training and testing parameters arises from its mixture-of-experts design: during training, all experts are active, whereas during inference, only the top-2 experts are selected, reducing the number of active parameters. RAIL-Primal incurs additional inference overhead due to its use of high-dimensional feature expansion through projection and regression layers. Meanwhile, RAIL-DUAL significantly increases memory usage by maintaining a memory bank of image features for all training samples across 1,100 classes. Notably, these methods experience linearly growing computational costs as the number of tasks increases. In contrast, our method minimizes training overhead by selectively retaining only ranks with high importance. Once training is complete, these updates are merged into the pre-trained weights, ensuring zero additional memory overhead during inference.
Figure 7: Ablations on different insertion locations and the impact of varying the initial rank.
Figure 8: Analyses (a) maximum pruning threshold and (b) ratios of dense training iterations.
4.4 Ablation Studies
Effects of insertion locations and number of initial ranks. We examine the impact of different insertion locations and initial ranks in our method, with results presented in Fig. 7. These findings are consistent with our analyses shown in Fig. 4. Our results indicate that when training is restricted to the vision encoder, transfer accuracy drops significantly, suggesting that the pre-trained capabilities are negatively impacted. Moreover, at specific locations, increasing the maximum rank improves adaptation ability while not always reducing transfer accuracy.
Interestingly, restricting updates to only the Attention modules can surpass our reported state-of-the-art Transfer metric in Table 1 and Table 2. However, this comes at the expense of reduced adaptation capabilities. To strike a better balance between adaptation and retention, we enable updates across all modules by default, ensuring an optimal trade-off between learning new tasks and retaining the model’s pre-trained capabilities.
Analyses of pruning threshold and dense-iteration ratio. To better understand the trade-off between downstream learning and forgetting, we control the number of adaptive ranks by varying the maximum threshold $\kappa_{\max}$ and the fraction of dense training iterations (Fig. 8). Gradually raising $\kappa_{\max}$ improves Last accuracy by pruning more useless or noisy ranks, whereas overly aggressive pruning limits task adaptation; a modest setting of 0.005 strikes a good balance. The dense-iteration ratio governs pruning onset: pruning from the start (ratio = 0) removes ranks before it can learn from the training data, while delaying pruning too long increases the risk of retaining useless or noisy ranks that do not get pruned eventually.
5 Conclusion
In this paper, we introduced CoDyRA, a novel approach to addressing continual learning challenges in vision-language models. By dynamically adjusting ranks based on their importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead. Extensive evaluations across various continual learning settings demonstrate that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and pre-trained retention.
Limitations and future works. Our approach merges updates into pre-trained weights without modeling inter-task correlations, mainly evaluated using pre-trained CLIP model. Future works could explore leveraging inter-task relationships and extending our method to other PTMs and settings.
Acknowledgments and Disclosure of Funding
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
Appendix A Additional Details of the Proposed Method
In Sec. 3.3 of the main paper, we introduce our method for learning dynamic rank-selective parameter updates to pre-trained weight matrices. We leverage a learnable vector to represent the importance of each rank, which is updated through a soft-thresholding operation. In this section, we provide more details to enhance clarity and understanding.
A.1 Full Derivation of Eq. (4)
For simplicity, we omit the notation for the training task $t$ in the following derivations. With a slight abuse of notation, we use subscript $t$ to represent $t$-th training iteration. At the $t$-th training iteration, the training loss with $\ell_{1}$ sparse regularization, as defined in the main paper, is:
$\mathcal{L}_{\text{train}}(\Delta_{t}):=\mathcal{L}_{\text{sup}}(\Delta_{t})+\lambda\sum_{m=1}^{M}\|\mathbf{w}^{m}_{t}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}$ is the supervised training loss, $\Delta_{t}$ is the all trainable parameters $\{\mathbf{w}^{m}_{t},\mathbf{B}^{m}_{t},\mathbf{A}^{m}_{t}\}_{m=1}^{M}$ at $t$-th iteration, $M$ is the total number of pre-trained weight matrices with LoRA of rank $r$, $\mathbf{w}^{m}_{t}$ represents the importance weights of the $m$-th LoRA module, and $\lambda>0$ controls the strength of the $\ell_{1}$ regularization.
We apply proximal gradient descent [4] to handle the non-differentiable $\ell_{1}$ regularization of the importance weights [12]. At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
$\mathbf{w}_{t+1}^{m}$ $\leftarrow\arg\min_{\mathbf{w}^{m}}\,\eta_{t}\cdot\lambda\|\mathbf{w}^{m}\|_{1}$
$\quad+\frac{1}{2}\|\mathbf{w}^{m}-(\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t}))\|_{2}^{2},$ (5)
where $\eta_{t}>0$ is the step-size of $t$-th iteration, and $\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})$ is the gradient of the supervised loss with respect to $\mathbf{w}^{m}$. This update balances the $\ell_{1}$ sparsity regularization term $\|\mathbf{w}^{m}\|_{1}$, with the quadratic penalty term $\frac{1}{2}\|\mathbf{w}^{m}-\hat{\mathbf{w}}^{m}_{t}\|_{2}^{2}$ , which ensures proximity to the gradient update if importance weights $\hat{\mathbf{w}}^{m}_{t}:=\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})$. This can be achieved through a soft-thresholding operator $\mathcal{T}$:
$\mathbf{w}_{t+1}^{m}\leftarrow\mathcal{T}_{\eta_{t}\cdot\lambda}(\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})),$ (6)
where the soft-thresholding operator $\mathcal{T}$ is defined as:
$\mathcal{T}_{\kappa}(x)=\mathbbm{1}(\lvert x\rvert>\kappa)\cdot(x+\mathrm{sign}(x)\cdot\kappa),$ (7)
such that the input to the operator $\mathcal{T}_{\kappa}$ is the elements in $\hat{\mathbf{w}}^{m}_{t}$, $\hat{\mathbf{w}}^{m}_{t}$ is the value of $\mathbf{w}^{m}_{t}$ after the gradient update using only $\mathcal{L}_{\text{sup}}(\Delta_{t})$, and $\kappa=\eta_{t}\cdot\lambda$ is the threshold. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met, and 0 otherwise, while $\mathrm{sign}(\cdot)$ denotes the sign (+/)(+/-)( + / - ) of the input. This concludes our derivation of the dynamic update for the importance weights as presented in the main paper Eq. (4).
Appendix B More Experimental Results and Ablations
Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper. We report experiment results averaged over 3 independent runs for reproducibility. Experiments were conducted on a single NVIDIA GeForce RTX 3090 GPU.
B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
The related continual learning methods focus on updating the representation with the backbone model (e.g., [53, 66, 61], and CoDyRA) or newly added representation alignment or prediction head [37, 58]. These two types of methods can be seen as orthogonal. We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in [58].
We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 5. Note that the Dual version of RAIL [58] stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 2 for a fair comparison.
RAIL is a regression-based method that applies an additional prediction head on top of the visual representations extracted by the frozen pre-trained vision encoder of the CLIP model. This approach is fully orthogonal to our method, as our focus is on continually incorporating new knowledge into the pre-trained model through dynamic low-rank parameter updates, and we can seamlessly integrate their techniques into our method. Because RAIL heavily relies on the frozen pre-trained model, the Transfer accuracy is limited by the zero-shot performance of the pre-trained model. In contrast, our approach gradually accumulates knowledge into the pre-trained model, enhancing its capability on unseen data and achieving improved Transfer accuracy. Moreover, by leveraging our continually enhanced pre-trained model, coupling our method with RAIL significantly outperforms the original RAIL, which uses a frozen pre-trained model, on the Average and Last metrics.
Table 5: Experimental results for integrating aggression-based adapters in the X-TAIL setting. Metrics for “Transfer”, “Average”, and “Last” are reported as averages across all datasets.
Transfer Average Last
RAIL-Primal 62.4 70.7 79.1
CoDyRA + RAIL-Primal 63.1 74.1 83.4
RAIL-Dual 62.4 71.9 82.4
CoDyRA + RAIL-Dual 63.1 74.2 83.6
B.2 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
In Sec. 4.3 of the main paper, Fig. 6 and Fig. 6 visualize and analyze the dynamic ranks assigned to each module. Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in [21]. The amplification factor is given by:
$\text{Amp}=\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{\text{T}}\mathbf{W}\mathbf{V}^{\text{T}}\|_{F}},$ (8)
where $\mathbf{W}$ is the pre-trained weight matrix, and $\Delta\mathbf{W}$ is the parameter change introduced by LoRA. For vanilla LoRA and CoDyRA, $\Delta\mathbf{W}$ is computed using Eq. (1) and Eq. (2), respectively. $\mathbf{U}$ and $\mathbf{V}$ represent the top $r$ singular vectors of $\text{SVD}(\Delta\mathbf{W})$, where $r$ is the number of interested ranks in analysis.
A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights [21]. We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig. 9.
The results show that, compared to vanilla LoRA, our method tends to concentrate parameter updates on a more specific subset of transformer modules while reducing updates to pre-trained weights less relevant to the current data. This demonstrates that our method more effectively identifies critical pre-trained weight matrices relevant to the current data and applies targeted dynamic parameter updates. Additionally, by reducing the strength of parameter updates for pre-trained weights less related to the current data, our method minimizes the impact on pre-trained knowledge and knowledge acquired from previous tasks compared to vanilla LoRA.
Note that this analysis differs from the visualizations of number of activated ranks in Fig. 6 and Fig. 6. The number of activated ranks reflects the quantity of significant ranks contributing to parameter updates for leaning each task. In this analysis, the amplification factor directly measures the parameter change to each pre-trained weight matrix.
(a) Aircraft - LoRA
(b) EuroSAT - LoRA
(c) Pets - LoRA
(d) Aircraft - CoDyRA
(e) EuroSAT - CoDyRA
(f) Pets - CoDyRA
Figure 9: Visualization of amplification factors when trained on (a, d) Aircraft, (b, e) EuroSAT, and (c, f) Pets, comparing (d, e, f) our method of dynamic sparse rank selective LoRA and (a, b, c) vanilla LoRA. It shows that the proposed method with adaptive rank selection can achieve more focused/concentrated updating and less interference.
B.3 More Visualizations and Analyses on Rank allocation results of different datasets
In Sec. 4.3, Fig. 6 and Fig. 6, we visualize and analyze the dynamic ranks assigned to each module for the Aircraft and Oxford Pets datasets under the X-TAIL experimental settings. Here, we extend these visualizations and analyses to additional datasets, as shown from Fig. 11 to Fig. 17.
The visualizations, along with statistical analyses of transformer modules and layers, reveal distinct rank allocation patterns across datasets. These findings suggest that the rank of parameter updates needed for achieving a good downstream improvements and knowledge retention is distinct across each kind of data.
Figure 10: Visualization and statistical analysis of rank activation on the Caltech101 dataset using our proposed method.
Figure 11: Visualization and statistical analysis of rank activation on the DTD dataset using our proposed method.
Figure 12: Visualization and statistical analysis of rank activation on the EuroSAT dataset using our proposed method.
Figure 13: Visualization and statistical analysis of rank activation on the Flowers dataset using our proposed method.
Figure 14: Visualization and statistical analysis of rank activation on the Food101 dataset using our proposed method.
Figure 15: Visualization and statistical analysis of rank activation on the MNIST dataset using our proposed method.
Figure 16: Visualization and statistical analysis of rank activation on the Stanford Cars dataset using our proposed method.
Figure 17: Visualization and statistical analysis of rank activation on the SUN397 dataset using our proposed method.
B.4 Evaluation with Other VLMs
To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP [28] on X-TAIL (Table 6). Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
Table 6: Averaged performance using BLIP.
Transfer Average Last
Zero-shot 51.03 46.45
[…]
Table 7: Performance of 16-shot PEFT on downstream tasks.
Acc. $\uparrow$ / Final Used Params. $\downarrow$
Caltech101 Cars OxfordPet
Zero-shot 88.40 64.70 89.00
LoRA ($r=16$) 96.43 / 4.4M 80.66 / 4.4M 92.94 / 4.4M
CoDyRA  ($r_{init}=16$) 97.20 / 3.95M 81.61 / 3.75M 94.44 / 3.85M
B.5 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
Leveraging our dynamic rank-adaptive parameter updates, our method naturally extends to Parameter-Efficient Fine-Tuning (PEFT) tasks. Table 7 presents 16-shot PEFT results, using the same hyperparameters as the main paper, showing that CoDyRA outperforms vanilla LoRA while reducing final trainable parameters by pruning ranks with zero importance. Notably, across multiple downstream tasks (Caltech101, Cars, and OxfordPet), CoDyRA achieves higher accuracy while using fewer parameters compared to LoRA with a fixed rank of 16.
These results highlight a more adaptive and efficient alternative to conventional low-rank tuning by automatically selecting the most relevant ranks based on training data. Crucially, by amplifying important ranks through learned higher importance weights, our approach significantly enhances downstream adaptation.
Broader Impacts
CoDyRA’s sparse, adaptive, importance-aware low-rank subspace learning significantly reduces compute and memory demands for continual learning, enabling on-device personalization in domains such as mobile health and education and empowering resource-constrained users and organizations. As a fundamental research effort, CodyRA does not address potential biases or hazards in training data, and its incremental updates could introduce privacy risks or stealthy backdoors. We therefore recommend transparent update logging, regular audits, and differential-privacy safeguards. Although more efficient than full fine-tuning, large-scale deployment still incurs energy costs; developers should monitor resource use, leverage low-power hardware, and explore carbon offsets.